Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b
(4,7); x+9y=4
I am completely stuck on this one. Any help wpuld be appreciated. Thank you.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

These are all the same.
Try doing it like this.
A line and a point example.
Write in standard form the eqation of a line that satisfies the given conditions. Perpendicular to 9x+3y=36, through (1,2)
-----------------
Find the slope of the line. Do that by putting the equation in slope-intercept form, y = mx + b. That means solve for y.
9x+3y = 36
3y= - 9x + 36
y = -3x + 13
The slope, m = -3
------------------
The slope of lines parallel have the same slope.
The slope of lines perpendicular is the negative inverse, m = +1/3
----------------
Use y = mx + b and the point (1,2) to find b.
2 = (1/3)*1 + b
b = 5/3
The equation is y = (1/3)x + 5/3 (slope-intercept form)
x - 3y = -5 (standard form)
------------------------
For further assistance, or to check your work, email me via the thank you note.

Answer by richwmiller(17219)

If you put a little effort into finding the right category, you would have been able to find many similar questions.
This is probably the most common problem submitted.
Did you look in your book or notes? You do have notes, don't you?